OpenStudy (anonymous):

Can someone please help me solve 3*6=-9+k?

Nnesha (nnesha):

what is 3 times 6

Nnesha (nnesha):

:)

OpenStudy (anonymous):

48

Nnesha (nnesha):

3 times 6 = 48 ???

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

is that wrong?

Nnesha (nnesha):

no use calculator :)

Nnesha (nnesha):

ye s that is wrong

Nnesha (nnesha):

are you kidding me

OpenStudy (anonymous):

6+6+6 = 18

Nnesha (nnesha):

ok so then calculator is lair never use that ok :)

OpenStudy (anonymous):

woops. Sorry I was supposed to put in 16, not 6. My mistake. The equation is 3*16=-9+k

Nnesha (nnesha):

ok so multply 16 times 3 is 48 = -9 + k now you have to move 9

Nnesha (nnesha):

move -9******

Nnesha (nnesha):

and 9 is subtracting on right side what is opposite of subtract

OpenStudy (anonymous):

addition

OpenStudy (kendricklamar2014):

This is how u solve the problem @Kirstenjla : Let's solve your equation step-by-step. (3)(16)=−9+k Step 1: Simplify both sides of the equation. 48=k−9

OpenStudy (kendricklamar2014):

Step 2: Flip the equation. k−9=48 Step 3: Add 9 to both sides. k−9+9=48+9 k=57

Nnesha (nnesha):

so add 9 both side

OpenStudy (kendricklamar2014):

Answer: k=57

OpenStudy (anonymous):

Thank you KendrickLamar2014.

Nnesha (nnesha):

9-9 should cancel out 48 + 9 = ?????????????????????

OpenStudy (anonymous):

57

OpenStudy (kendricklamar2014):

You'r Welcome

Nnesha (nnesha):

well ok lets suppose that i'm not sure that 57 is right or not so we should check by plugging back into the equation like  yep both side are right os yay!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! you got right answer